A shipment of 30 inexpensive digital watches, including 10 that are defective, is sent to a department store. The receiving department selects 10 at random for testing and rejects the whole shipment if 1 or more in the sample are found defective. What is the probability that the shipment will be rejected? 


The probability the shipment will be rejected is _______ . 

(Simplify your answer. Type an integer or decimal rounded to two decimal places as needed.)
Answer #1

30 Shipment of Including watches 10 inexpensive digital Wat Ches that are defective testing randomlig taken 10 for No. of wat
Answer #2
[REMOVED]
Answer #3

SOLUTION :


In the lot of 30 watches 10 are defective.

So, 20 out of 30 are not defective.


P ( No defective in the randomly  chosen 10 out of 30 ) (without replacement)

= 20/30 * 19/29 * …… and so on ….. * 11/21

= (20! / 10!) / (30! / 20!)

= (20! * 20!) / (30! * 10!)

= 0.0061493 


So, 


P( rejecting the lot)

= P(1 or more defective in 10 chosen out of 30)

= 1 - P (No defective in 10 chosen out of 30)

= 1 - 0.0061493

= 0.9938507 

= 0.99 (to two decimal places) 


So, Probability of rejecting the lot is 0.99  (ANSWER).


answered by: Tulsiram Garg